Two point charges are on the y axis. A 4.90-µC charge is located at y = 1.25 cm, and a -2.3-µC charge is located at y = −1.80 cm.

(a) Find the total electric potential at the origin.
_____________ V

(b) Find the total electric potential at the point whose coordinates are (1.50 cm, 0).
_____________V
Answer #1

Potential at origin, V = k*q1/r1 + k*q2/r2

= 9*10^9*(4.9*10^-6)/0.0125 + 9*10^9*(-2.3*10^-6)/0.018

= 2.38*10^6 V

b)

Potential at the point ( 1.5cm , 0 )

V = k*q1/r1 + k*q2/r2

= 9*10^9*(4.9*10^-6)/sqrt(0.0125^2 + 0.015^2) + 9*10^9*(-2.3*10^-6)/sqrt(0.018^2 + 0.015^2)

= 1.38*10^6 V

> what if for b it's (0, 1.25)

Preston Cottrell Sun, Mar 13, 2022 8:49 PM

> do the Pythagorean Theorem and solve for c which is the distance from the charge to the point (0, 1.25)

Preston Cottrell Sun, Mar 13, 2022 9:05 PM